# Ringleader ASGD: The First Asynchronous SGD with Optimal Time Complexity under Data Heterogeneity

### v2

@@ Appendix B Arbitrarily Changing Computation Times
B.1 Universal Computation Model
@@
B.2 Toward an Optimal Method
@@ B.3 An Optimal Method
Appendix C Auxiliary Lemmas
@@ C.2 Variance Term
$\bar{g}^{k}=\frac{1}{n}\sum_{i=1}^{n}\bar{g}_{i}^{k}=\frac{1}{n}\sum_{i=1}^{n}g_{i}^{k,j}=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}\left(x^{k-\delta_{i}^{k}};\xi_{i}^{k-\delta_{i}^{k},j}\right)~.$
@@
The inequalities follow from the Cauchy-Schwarz inequality, LL–smoothness of $f$, the triangle inequality, Young’s inequality, Section C.2, and finally $(k\bmod n)\leq n-1<n$.
@@
where in the last step we used Section C.2.
@@
In the last inequality, we used Jensen’s inequality and Section C.2.
@@
Appendix D Time Complexity of IA2SGD
@@
Appendix E Improved Malenia SGD
Malenia SGD has the following iteration complexity (Tyurin & Richtárik, 2024)
@@ Appendix E Improved Malenia SGD
for all iterations $k$. In the original Malenia SGD analysis (Tyurin & Richtárik, 2024), this bound follows from the condition in (3), which fixes the same value of $S$ across all iterations.
@@
Finally, note that this improvement is only valid in the fixed-time regime. In the setting with arbitrarily varying computation times, the same optimization cannot be applied, for the same reasons discussed for Ringleader ASGD in Appendix B.

### v3

@@ Appendix C Arbitrarily Changing Computation Times
C.1 Universal Computation Model
@@
C.2 Toward an Optimal Method
@@ C.3 An Optimal Method
Appendix D Auxiliary Lemmas
@@ D.2 Variance Term
$\bar{g}^{k}=\frac{1}{n}\sum_{i=1}^{n}\bar{g}_{i}^{k}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{b_{i}^{k}}\sum_{j=1}^{b_{i}^{k}}g_{i}^{k,j}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{b_{i}^{k}}\sum_{j=1}^{b_{i}^{k}}\nabla f_{i}\left(x^{k-\delta_{i}^{k}};\xi_{i}^{k-\delta_{i}^{k},j}\right)~.$
@@
The inequalities follow from the Cauchy-Schwarz inequality, LL–smoothness of $f$, the triangle inequality, Young’s inequality, Section D.2, and finally $(k\bmod n)\leq n-1<n$.
@@
where in the last step we used Section D.2.
@@
In the last inequality, we used Jensen’s inequality and Section D.2.
@@
Appendix E Time Complexity of IA2SGD
@@
Appendix F Improved Malenia SGD
Malenia SGD has the following iteration complexity (Tyurin and Richtárik, 2024b)
@@ Appendix F Improved Malenia SGD
for all iterations $k$. In the original Malenia SGD analysis (Tyurin and Richtárik, 2024b), this bound follows from the condition in (3), which fixes the same value of $S$ across all iterations.
@@
Finally, note that this improvement is only valid in the fixed-time regime. In the setting with arbitrarily varying computation times, the same optimization cannot be applied, for the same reasons discussed for Ringleader ASGD in Appendix C.